the length of a rectangle is 6 centimeters and its area is (6x+18) square centimeters . what is an expression for the width.
Answer by macston(5194)

.
L=length; A=area; W=width
.
A=LW
A/L=W
(6x+18)/6=W
x+3=W
ANSWER: The width=x+3.